mastering physics 22.6

uploaded imageThe cube in thefigure has sides of length L = 10.0 cm . The electric field is uniform, has a magnitude E = 4.00 * 10^3 N/C, and is parallel to the xy-plane at an angle of 36.9degrees measured from the + x axis toward the + y axis.

a.What is the electric flux through the cube face S_1?

b. What is the electric flux through the cube face S_2?

c. What is the electric flux through the cube face S_3?

d. What is the electric flux through the cube face S_4?

e. What is the electric flux through the cube face S_5?

f. What is the electric flux through the cube face S_6?

g. What is the total electric flux through all faces of the cube?
Answer #1

We have electric fluxΦ = E A cosθ

given E = 4*103N/CArea A = (10*10-2m)(10*10-2m) =0.01m2θ is the angle made by the electric field with the normalto the surface.A) For the surface S1normal to the surface is -yaxis.here the electric field makes an angle 36.9owiththe positive x - axis. therefore angle made by the electric fieldwith the y-axis is90o-36.9o=53.1otherefore the angle made by the electric field with the normalto the surface S1is 180o+ 53.1o= 233.1oThenΦ1=(4*103N/C)(0.01m2)(cos233.1o)=-24Nm2/CB) For the surface S1normal to the surface isz axis.as the elcetric field is parallel to the XY plane thereforetheangle made by the electric field with the z-axis is90otherefore the angle made by the electric field with the normalto the surface S2is 90oThenΦ2=(4*103N/C)(0.01m2)(cos90o)=0Nm2/CC) For the surface S3normal to the surface isy axis.here the electric field makes an angle 36.9owiththe positive x - axis. therefore angle made by the electric fieldwith they-axis is90o-36.9o=53.1otherefore the angle made by the electric field with the normalto the surface S3is 53.1oThenΦ3=(4*103N/C)(0.01m2)(cos53.1o)=24Nm2/CD)For the surface S4normal to thesurface is z axis.as the elcetric field is parallel to the XY plane thereforetheangle made by the electric field with the z-axis is90otherefore the angle made by the electric field with the normalto the surface S4is 90oThenΦ4=(4*103N/C)(0.01m2)(cos90o)=0Nm2/CE) For the surface S5normal to the surface isx axis.here the electric field makes an angle 36.9owiththe positive x - axis.therefore the angle made by theelectricfield with the normal to the surface S5is36.9oThenΦ5=(4*103N/C)(0.01m2)(cos36.9o)=32Nm2/CF) For the surface S6normal to the surface is-x axis.here the electric field makes an angle 36.9owiththe positive x - axis.therefore the angle made bytheelectric field with the normal to the surface S6is180o+36.9o=216.9oThenΦ6=(4*103N/C)(0.01m2)(cos216.9o)=-32Nm2/CG)Total electric flux isΦ=Φ123456= 0Nm2/C
answered by: Quin